OpenStudy (anonymous):

The equation of line EF is y = 2x + 1. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0, 2). y = 2x − 4 y = 2x + 2 y = negative 1 over 2x − 4 y = 1 over 2x + 2

OpenStudy (anonymous):

@Agl202

OpenStudy (lexluth0r):

You would have to use point-slope form, which is y-y1=m(x-x1), where (x1,y1) is a given point and m is the slope.

OpenStudy (lexluth0r):

You can plug into the equation like this: y-2=2(x-0) because (0,2) is the given equation, and the slope is 2 because to be parallel tot he first line the slope has to match.

OpenStudy (lexluth0r):

to the*

OpenStudy (lexluth0r):

Then you can simplify the equation to slope-intercept form

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

so then i get??

OpenStudy (lexluth0r):

It should simplifiy to y=2x+2

OpenStudy (anonymous):

Ohh okay thank you sooo much!!

OpenStudy (lexluth0r):

NP! Glad to help C;